Monthly Report IR Demo FEL Upgrade Project
September 2000

Project Highlights

September 2000 was the fourth month of activity for the IR Demo FEL Upgrade Project.  Design and initial fabrication activities for the IR Demo continued as noted below.  The project was given a favorable review
during the annual review of Jefferson Lab's Science and Technology held by DOE and SURA on Sept. 25-27th.  Cost and schedule performance for the Upgrade Project are given in the appended "Performance Assessment

Project Cost Performance

The project budget for the period June 1, 2000 to Sept. 30, 2001 is $9,029k. The project through the month of September has a total of $1,193k of performance scheduled (assuming the project started at the originally planned start date of April 1, 2000).  The work performed through the end of September was $1,427k, which is 16% complete vs. 13% scheduled.  The actual cost accrued through August totals $1,047k.  This results in a schedule variance of +$235k and a cost variance of +$380k.  The favorable cost and schedule variance continues because of the large quantity of RF work (WBS 6) that was completed ahead of schedule and under cost.  With the current month's accrual, the RF work complete to date is $718k or 75% complete.

Management Highlights:

We began planning with the FEL team and with our Navy contract monitor (J. Albertine), for the agenda for the next 6-month review on the FEL Upgrade project, which is scheduled for Oct. 23rd at Jefferson Lab.

The FEL Project received an excellent review at the SURA/DOE review of Jefferson Lab’s Science and technology program on Sept. 25-27.  The review panel thought that the Upgrade project was the right and proper
next step after the kilowatt demo. The committee recognized our challenge to energize the user community behind scientific applications of the present and upgraded FEL.  They thought the FEL had already demonstrated and had a strong potential to further enrich and diversify the Jefferson lab mission.

A presentation on biological applications of the FEL was given to the Virginia Joint (House/Senate) Subcommittee on Science and Technology.  The group is investigating ways the state legislature can encourage the development of biotechnology in Virginia.  The role of the FEL as an attraction to the “superstars” of the field was discussed.  Steve Benson and Jim Boyce represented the FEL team and gave presentations at the Opto Southeast meeting held in Charlotte on Sept. 18-19.  Michelle Shinn represented the team at a DOD Applications of Lasers
meeting, which was held Sept. 25-27 at the Pennsylvania State University.  Of the 100 people attending, a good cross-section of industry and government was represented.  A small group visited Fishkill, NY, to investigate the compatibility of the IR Upgrade design with higher energy (x-ray) photon production.  Physical interferences, control system status, and documentation were all reviewed.  No major hurdles were seen.

WBS 3.0 (Beam Physics):

The following is summary for the weeks of 9/11 and 9/19:
Longitudinal phase space dynamics for the upgrade were investigated and scenarios for nominal operation were established. In addition, a configuration leading to ultra-short bunches (~10-20 fsec rms) was developed.  This work is documented in JLAB-TN-00-020, available on line at
Momentum acceptance measurements were made on the driver; the latest (9/22) indicated the full momentum aperture was only of order 5%, in contrast to the 6+% indicated by FEL operation.  Fascinating nonlinear effects (such as T346 induced vertical dispersion) were observed and corrected by orbit ironing; additional work involved implementation of the multi- (quad and sextupole) family compaction and dispersion management schemes considered necessary in the upgrade.  Initial work on
UV bypass concepts yielded four candidate endloop designs, 2 based on a stair-case (reverse bending with dispersion modulation) geometry and two based on a single step (dispersion modulation only) geometry.  Analysis of each in a complete recirculator layout will start shortly.  Beam dynamics and engineering personnel initiated discussion of the energy recovery dump and developed layout options for both upgraded Demo and large aperture "garbage can" style dumps.  Once established, these
footprints will provide guidance for the beam optics design of, and instrumentation for, this region.

We began to organize our efforts needed to consider changes in the electrostatics of gun electrode geometry.  The upgraded gun may allow a slightly larger and different shape cathode, which can help the high field performance.  We are going to start by resuscitating the original electrostatic model done by D. Engwall and work from there to evaluate potential modifications.  B. Yunn will be spearheading this task.  Work continued on modifications of the lattice that will allow the UV chicane/wiggler extension to be added to the baseline upgrade lattice.  This modification will be a component of the planned UV program, which
has been proposed to AFOSR for start-up funding in FY01.

WBS 4.0 (Injector):

Light Box preliminary design issues are being discussed.  Final concepts discussions are planned.  The field emission test system has been running with a 9um polished N2 implanted SST cathode and a 1um polished SST anode at a 5 mm gap with excellent results.  The gap has held off 24 MV/m with a very small dark current (145 pA).  The conditioned gap at 21MV/m is drawing 90 pA.  These results present an excellent candidate material for the upgrade gun electrode.

WBS 5 (SRF):

In the cavity area, we tested the new vacuum seal arrangement for the 7-cell cavities that involves the use of Al-Mg diamond-shaped gaskets with Nb-Ti flanges.  The flanges have successfully passed ambient temperature leak tests.  Cold tests will follow.  Nb cells for the first four cavities have been pressed and will now move to frequency measurement and weld-up.  Material is on hand for the remaining 4 cavities.  We received the final proposal from the vendor that will be manufacturing the warm-to-cold waveguide sections.  The proposal is under evaluation.

Had an initial vendor visit to start vacuum vessel and space frame fabrication. No outstanding cost, technical, or schedule issues. Vendor is well prepared for contract.  Waveguide and end can procurements continue to make progress.  Signed off top hat design drawings.  Heat shield design package is almost complete.  We had discussions with the SRF team (J. Mammosser) and our collaborator from KEK (K. Saito) about KEK’s proposal to treat cavities by electropolishing.  Currently, we have sent one test 7-cell cavity to KEK and a second one will be sent in the next month or so.  As reported at the LINAC conference in August, KEK has produced by electropolishing several 9-cell cavities with gradients greater than 30MV/m.  KEK will make a formal proposal to Jefferson Lab to treat all eight, 7-cell cavities for the FEL upgrade cryomodule in exchange for our performance test data.

RF Systems (WBS 6):

The installation of the RF hardware for the third cryomodule position (Zone 4) is essentially 95% complete.  Only the klystrons remain to be installed and then testing of the Zone 4 RF system can commence.  For the second cryomodule position (Zone 3), the low level racks are complete and the HPA is about 50% complete. Procurements have been written for the HPA HVPS, the circulators and the klystrons.  We signed off on the technical specification for the HVPS for the HPA for the 100kW RF system in the injector (Zone 2).  And it was sent to procurement to begin the order process.

We are working with Glassman HV to provide a slave unit for the Gun HVPS to achieve 10 ma beam current.

Instrumentation (WBS 8):

Preparations continued in getting the contracts on the street for the 3" BPMs (44) and for the 3" shielded beam viewers (36).  The prices are >$50 K for BPMs, and > $150 K for the viewers. In addition to having the jobs bid locally packets will be sent out to qualified vendors that Beam Instrumentation people from BNL, APS, Fermi, and LANL have used.  This will get us the best product & price.  These contracts given time to bid and pass through procurement should be awarded in 3 to 4 weeks.
Improvements continue to be made on electron beam and FEL diagnostics; the new sample/hold & integrator system is tested and operational.  The FEL main control menu has been gone over to remove obsolete and redundant control screens.  This will ease the transition to having CEBAF operators controlling the various systems.  Also progress has been made on user lab patch and network systems.  The design for the 32 channel VME based sample and hold board is complete and in EECAD for drawings and board layout. We hope to have a prototype to test during the Oct. run. The optical BPM design is proceeding well this should be ready for a prototype surface mount board layout by next week.  The user patch panels and user LAN connections are being worked on and will be complete by the Oct. run. Also, the 32 channel pico-motor multiplexer is
being finished up to allow control of turning mirrors in each of the user labs.

We accomplished significant work on the development and deployment of LiTaO3 pyroelectric sensors. These will be used as both power monitors and optical beam position monitors.  Five 10 mm square detectors are being prepared for power monitoring at various brewster window ports.  These are position independent as well as wavelength independent for maintaining a fixed average power to a given experiment whether the beam is pulsed (to 10 microseconds) or CW.  The electronics have been
somewhat challenging to provide sufficient gain and response time with limiting the noise, rise times on order of a few microseconds is desired.  These outputs run to the sample & hold card then to the A/D.  The Optical Beam Position Monitors (OBPM) use the same type crystal configured as four quadrants of 1mm X 1mm each.  The new 32 channel VME
sample & hold card layout is nearly complete, the initial boards will arrive late next week and will be operational ~10/10/00 just prior to the start of the run.  Progress continues to be made on providing pico-motor control for two turning mirrors in each of the six optical labs.  A chassis, which multiplexes one pico-motor channel to 32 individual outputs, was completed this week.  The VME control card is the FEL standard general purpose I/O.

Beam Transport System (WBS 9):

Details of the steel yoke parts of the GX Prototype are complete.  The 3D model the of the GX dipole, using the optimum bump on the pole edge was constructed and ran to completion.  The results are being processed to assess the improvement.
Poles:  L&L Tool machined the first of four poles.  We will measure it for conformance to the ideal shape. The Survey & Alignment Group performed a detailed CMM inspection of the First Article quadrant.  The inspection showed that most of the milled surfaces were within tolerance but that there was a small systematic offset from pole to back leg and most of the holes were out of tolerance.  Dave McCay and Jeff Karn are working the problem with the vendor.  J. Dail attached the cooling plates to the quadrupole coil set for the prototype.

Other Transport Hardware:
 Differential Pumping stations are being modified to accommodate a three-inch aperture. We made two layouts integrating the wide-angle beam dump and its required shielding into the machine so Dave Douglas could continue the design iteration.

Wiggler (WBS 10):

Manifold fabrication continued.  We measured the pole tip gaps.  They have about a 0.002 inch gradient and the center gaps were larger than the ends by several thousandths.  We are working on improving those readings and making them reproducible.
An assembly procedure guaranteeing better gap uniformity is being developed.

We formulated a plan to see if we could bring the pole tips into conformance with the gap specifications by simple assembly technique. If the plan is unsuccessful or not repeatable for several re assemblies, we may be forced to use the clamps and spacer bars that we thought we could abandon.  An insulator plate and its mounts for the main leads of the NG Wiggler were missing from the drawing package and will have to be drawn and fabricated.  Copies of the drawings and specifications were
made for the dispersion section.  In discussions with BNL we found out that we cannot use their magnetic measurement apparatus to characterize our wiggler sections because the maximum length is limited to 2m with the BNL apparatus.  We are investigating options using ANL or commercial equipment.

The top assembly drawing of the Optical Klystron was completed, reviewed and signed-off.  The Dispersion Section package is out for bid.

Optics (WBS 11):

Conceptual designs of the Optical beam Position Monitoring (OBPM) system were reviewed and approved.  The system will be tested in the existing Optical Transport System.  Monitoring detectors have been received and position controllers of several types will be ordered, tested, and evaluated as to compatibility and reliability.  Pyroelectric array diagnostics for spectral measurements were installed in the OCR for characterization.  Initial use will be manual with the full software suite added when functionality has been confirmed.  A pyroelectric detector with chopper was added to the OTS as a real time macropulse
power diagnostic when married with a new sample and hold circuit.  This will be tested on the next laser run.  The Modemaster system was realigned and checked out in preparation for more detailed studied of mirror heating distortion and its effect on the FEL lasing process.  A designer is preparing drawings of the planned mirror test stand which will be used to characterize performance of the controllable curvature mirror. We will decide soon which system to move.

We spent considerable time on the design of the new optical cavity which will be presented to the review committee at the Oct. 23 project review.  We recalculated parameters for a ~ 32m near-concentric cavity and began rewriting requirements document, based on the new parameters.

Beam Operations

The FEL was operated for one week in September for Gunter Luepke from the Dept. of Applied Science at the College of William and Mary.  Gunter and his team are continuing their studies of the dynamics of hydrogen vacancy complexes in silicon (their first results were recently published in Phys. Rev. Lett 85,1452, 2000).  His preliminary results from this week’s work appear equally as exciting and worthy of publication.  No further user runs are scheduled until the October run.  As usual, our current user requests for the Fall run exceed the run period time allotment by a significant margin.

We operated the FEL this week two days this month to make beam transport measurements.  This will help determine the energy aperture of the upgrade.

Privacy and Security Notice